### v1

@@ 4.1 Computing effective energy utilization $\eta$
2014 28 2.76 96.875 55 0.017848 0.63786
2018 16 1.64 23.4375 30 0.029776 0.438596
2022 7 0.75 10 6.25 0.044118 0.615385
Table 1: Energy utilization trends over time for conventional and $\pi^{2}$ architectures across various technology nodes. Energy efficiency ($E$), measured in picojoules per bit (pJ/bit), is used as the key performance metric. The temporal evolution of switch technology is derived from data reported in [53]. Projected compute energy efficiency values are interpolated from datasets presented in [62]. Switch energy efficiency figures are sourced from [63], [52], and [64], while interconnect energy efficiency data are drawn from [65], [53], and [66].
In large-scale conventional (neuromorphic) architectures with passive interconnects, multiple compute units exchange information via Ethernet-based protocols and switches, as illustrated in [12]. In contrast, large-scale $\pi^{2}$-NN systems can be realized using a single high-capacity Ethernet switch that supports millions of neurons. For conventional architectures, the energy utilization metric, denoted as Conv-$\eta$, is defined as the ratio of compute energy to the total system energy, comprising compute, switch, and interconnect energy. Given that switch and interconnect components dominate the energy budget, Conv-$\eta$ approaches zero, indicating poor effective utilization. In $\pi^{2}$ architecture, the utilization metric $\pi^{2}$-$\eta$ is defined as the ratio of switch energy to the total energy consumed by the switch and interconnect. This formulation reflects the integrated nature of computation and communication in $\pi^{2}$ systems, where the switch itself performs both functions efficiently. If the interconnect delay is also taken into consideration while computing, then the utilization can be pushed to 1.
@@ 4.2 $\pi^{2}$ primitives and standard neural network operations
$\pi^{2}_{K}$ neuron is a variant of the TEMP neuron model: Let $\mathcal{U}=\{t_{1},t_{2},\ldots,t_{d}\}$ denote the set of input event arrival times. For simplicity, we assume that $t_{1}\leq t_{2}\leq\cdots\leq t_{d}$ For a fixed parameter $\gamma\geq 0$, the TEMP neuron generates a spike/event at time T when the following condition is satisfied [23]:
@@
$\pi^{2}_{K}$ neurons can be interpreted as mimicking inner-product computation on their inputs: At the core of the TEMP neuron model lies margin propagation (MP), a piecewise-linear approximate computing technique introduced in [67, 29, 23]. In [29], MP was employed to approximate inner products between vectors, and the error characteristics of this approximation were analyzed and quantified. Following a similar reasoning, the TEMP neuron model can also be interpreted as an approximation to the inner product. Let $\mathbf{w},\mathbf{x}\in\mathbb{R}^{D}$ be $D$-dimensional real vectors. For convenience, we define the sets
@@
where $TEMP(\cdot,\gamma)$ denotes the operator defined in Eq. 10. Eq. 13 highlights the differential formulation inherited from MP to approximate inner product. Since we have established the equivalence between TEMP and the proposed $\pi^{2}_{K}$ neuron, the same relationship can be expressed in terms of $\pi^{2}_{K}$ as
@@ 4.3 Training of $\pi^{2}$ neural networks on GPU
We adopt a knowledge transfer-based training strategy for the $\pi^{2}$-NN, wherein a conventional MAC-based neural network (Teacher network) is first trained to achieve baseline accuracy on a given task. The trained model’s parameters are then used to initialize the $\pi^{2}$-NN, which is subsequently fine-tuned to match the performance of the teacher network. For all software-based simulations, we refer to the conventional MAC-based network as the teacher network and its corresponding $\pi^{2}$-NN implementation as the student network. For training the student network, we employ a distillation loss that combines the conventional cross-entropy loss with a temperature-scaled Kullback-Leibler (KL) divergence between the soft output distributions of the teacher and student networks. At the input layer, to convert the pixel values (X) to differential spike timing (T), as specified in [23], we use the following strategy.
@@
Training ResNet-9 architecture on the CIFAR-10 dataset: The CIFAR-10 dataset, comprising 50,000 training and 10,000 test images across 10 classes, was preprocessed by applying standard data augmentation techniques commonly used in image classification tasks. Each training image was randomly cropped to $32\times 32$ pixels with a padding of 4 using reflection padding, followed by random horizontal flipping. The pixel values were then normalized using the dataset-specific channel-wise mean and standard deviation values: $(\mu_{R},\mu_{G},\mu_{B})=(0.4914,0.4822,0.4465)$ and $(\sigma_{R},\sigma_{G},\sigma_{B})=(0.2470,0.2435,0.2616)$. The teacher network was trained using the Adam optimizer with a learning rate of $1\times 10^{-2}$, a weight decay of $1\times 10^{-4}$, and a OneCycleLR scheduler to achieve the baseline performance. To train the $\pi^{2}$-NN student model on CIFAR-10, we adopt a knowledge distillation strategy, initializing the network using a pre-trained MAC-based ResNet9 teacher model. The K values corresponding to the 9 layers of the network were as follows:$K=[16,50,35,35,60,70,80,200,50]$. The $\pi^{2}$ network was trained using the Adam optimizer with a learning rate of $1\times 10^{-2}$, a weight decay of $1\times 10^{-8}$, and a OneCycleLR scheduler. Training was conducted over 60 epochs using a batch size of 16.
Training the ResNet-9 architecture on the CIFAR-100 dataset: We train the teacher ResNet-9 network on the CIFAR-100 dataset, comprising 50,000 training and 10,000 test images across 100 classes. Image preprocessing includes random cropping with 4-pixel padding, horizontal flipping, and normalization using channel-wise mean and standard deviation values of (0.5071, 0.4867, 0.4408) and (0.2675, 0.2565, 0.2761), respectively. The model is trained using the Adam optimizer with a learning rate of $1\times 10^{-2}$ and a weight decay of $1\times 10^{-4}$ for 50 epochs. A OneCycleLR scheduler is employed to modulate the learning rate with a three-phase annealing schedule. The K values used to train the $\pi^{2}$-NN student model on CIFAR-100 were [10, 50, 55, 65, 100, 100, 80, 200, 100]. The network was trained using the Adam optimizer with a learning rate of $1\times 10^{-2}$, a weight decay of $1\times 10^{-8}$, and a OneCycleLR scheduler. Training was conducted over 60 epochs using a batch size of 16.
@@ 4.4 OMNET++ implementation of the MNIST $\pi^{2}$ network
The nodes/neurons are simulated as TSN devices, compatible with the traffic-shaping protocols. The synaptic delays are quantized and encoded within the Priority Code Point (PCP) fields of the incoming Ethernet frames. The OMNeT++ simulator supports a three-bit Priority Code Point (PCP) that is part of the VLAN tag [27]. So, the synaptic delays are quantized to 3 bits. The input events are first routed to an Ethernet switch implementing the $\pi^{2}$-ATS protocol. The TET values are computed based on the PCP value of the frame (synaptic delay), and the events are buffered until their TET is reached. The delayed events are then transmitted to another Ethernet switch implementing the $\pi^{2}$-CBS protocol to simulate the $\pi^{2}_{K}$ dynamics. The implementation details are presented in the supplementary section.
@@ 4.6 Calibrating for finite-precision, jitter and packet drops
We implement two types of hardware constraints: (i) spiking time jitter and (ii) weight quantization, and (iii) random packet drops. To incorporate weight quantization (Fig. 7A), we employ quantization-aware training (QAT), allowing weights to be represented with reduced bit precision during inference. Specifically, we fine-tune the network for 10 epochs using quantized weights, starting from a pretrained model. To incorporate spiking time jitter (Fig. 7C), a random value of a Gaussian noise of a given standard deviation is added to each spiking time of the TTFS-network inputs and the outputs of each layer, similar to [30, 68]. We fine-tune the network for 10 epochs using jitter and quantized weights (8 bits and 4 bits), starting from the pre-trained model. For packet drops (Fig. 7B), we stochastically suppress events within a layer with a preset drop probability and assess the resulting performance; this condition is likewise fine-tuned for 10 epochs from the pretrained initialization. Across all robustness experiments, we use an initial learning rate of 1e-3 and keep all other hyperparameters identical to the original training configuration.
@@ 4.7 Computational complexity of $\pi^{2}$ networks
where $C_{m}$ and $C_{a}$ represent the computational complexity of performing a multiplication and addition operation. For an n x n $\pi^{2}$ network following the $K^{th}$ neuron paradigm, the number of addition operations performed at every node of the output layer depends on the value of K. The delaying, time sorting, and event dropping operations are inherently implemented using interconnects. The number of operations at the neuron level is
@@
In the hardware implementation of conventional networks, the interconnects do the work of transmitting data to and from memory, external interfaces, communicating modules, and between different submodules within the hardware substrate. The communication complexities and memory accesses are usually ignored while computing the computational complexity in conventional networks [69]. Accordingly, in both cases, we have only considered the complexity of neuron-level computations. To emphasize the hardware efficiency of this approach, it is noteworthy that a 16-bit full adder typically requires approximately 250–500 transistors [70], while a 16-bit multiplier demands over 4,000 transistors [71]. Furthermore, the energy cost of a floating-point multiplication can be up to four times that of a floating-point addition [72], underscoring the significant hardware and energy advantages of the $\pi^{2}_{K}$ neuron.
@@ Acknowledgements
The authors would like to thank Haresh Dagale (Principal Research Scientist, Department of Electronic Systems Engineering, IISc) for his insights on communication network protocols. The authors would like to also acknowledge the joint IISc-WashU Memorandum of Understanding, to facilitate the collaboration between the two institutions. M.S.R, C.S.T and S.C. participated in the Neuromorphic Integrated Circuits workgroup at the Telluride Neuromorphic and Cognition Engineering workshop in 2024, Colorado, USA, which was supported by the National Science Foundation grant ECCS:2332166. M.S.R, C.S.T and S.C. also acknowledge the 2025 Bangalore Neuromorphic Engineering workshop (BNEW 2025) where some of the $\pi^{2}$-NN discussions occurred. S.C. would like to acknowledge the WashU McDonnell International Scholars Academy for supporting travel to BNEW 2025 at IISc. C.S.T would like to acknowledge the Pratiksha Trust for funding support.
@@
M.S.R, C.S.T and S.C. formulated the concept of $\pi^{2}$-NN and the use of interconnect primitives. M.S.R performed all the simulations and benchmarking. C.B. contributed to improve the training of $\pi^{2}$-NN. All authors/co-authors contributed to the proofreading and writing of the manuscript.
@@
The specific Python codes used for training $\pi^{2}$ networks are available at https://github.com/NeuRonICS-Lab/Processing-in-Interconnect-Codes Omnet++ codes will be available on request.
@@
Supplementary Information
@@ 1 A simple $\pi^{2}$ inspired point-to-point (P2P) processor architecture
Figure 9: A) $\pi^{2}$ inspired P2P architecture. B) Roofline model of the proposed architecture
We propose a point-to-point $\pi^{2}$ network architecture (Fig.9A) where every pair of neurons is connected through a path with a specific routing delay. The synaptic weights in a conventional neural network correspond to interconnect delays here. The interconnect length—or the specific routing path taken by the data—is determined by the trained synaptic weight magnitudes and is pre-defined. Thus, whenever an input neuron sends out data, it gets delayed automatically before it reaches its destination. Also, the shared interconnect bandwidth to the Rx neuron can only accommodate K events. Since the events arrive in a time-sorted manner, the first K events will inherently be accommodated at the Rx node, and the remaining events will be dropped. This approach effectively integrates processing and memory within the interconnect itself, as it inherently delays, sorts, and drops events, eliminating the need for explicit synaptic weight storage. We assume the neurons possess sufficient local memory and compute to accumulate membrane potential values and generate an event at T (output spike time).
@@
$Ops/s=\frac{Ln^{2}+n+n(Ln-K)}{T}$ (21)
As can be observed, only increasing the number of interconnects in this architecture, or the bandwidth, can delay, drop, and transmit more events, thus improving the peak throughput of this processor architecture. Additionally, there is no access to any external memory, so the roofline model remains parallel to the X-axis, as shown in Fig. 9B. The latency of computation (T) will likely be further reduced, as the Rx neurons are likely to generate events faster (as they receive more data), which is bound to improve the throughput even further.
@@
If the interconnect between the Tx and Rx nodes has a bandwidth BW of L event/T(s), then the power consumed (W) in T seconds is:
$P_{useful}=\frac{Ln^{2}E_{t}+nE_{c}}{T}$ (22)
@@
$P_{lost}=\frac{E_{p}n(F-L)_{+}}{T}$ (23)
When interconnects contribute to both computation and transmission, the transmission energy becomes an integral part of the system’s operational cost rather than an overhead. The energy/power utilization of the system can be computed to be:
$\eta=1-P_{lost}/(P_{useful}+P_{lost})$ (24)

### v3

@@ 4.1 Computing effective energy utilization $\eta$
2014 28 2.76 96.875 55 0.017848 0.59
2018 16 1.64 23.4375 30 0.029776 0.71
2022 7 0.75 10 6.25 0.044118 0.61
Table 1: Energy utilization trends over time for conventional and $\pi^{2}$ architectures across various technology nodes. Energy efficiency ($E$), measured in picojoules per bit (pJ/bit), is used as the key performance metric. The temporal evolution of switch technology is derived from data reported in [cpo]. Projected compute energy efficiency values are interpolated from datasets presented in [compute_energy]. Switch energy efficiency figures are sourced from [marvel], [25.6], and [edgecore_switch], while interconnect energy efficiency data are drawn from [qsfp], [cpo], and [broadcommcpo].
In large-scale conventional (neuromorphic) architectures with passive interconnects, multiple compute units exchange information via Ethernet-based protocols and switches, as illustrated in [truenorth_large_scale]. In contrast, large-scale $\pi^{2}$-NN systems can be realized using a single high-capacity Ethernet switch that supports millions of neurons. For conventional architectures, the energy utilization metric, denoted as Conv-$\eta$, is defined as the ratio of compute energy to the total system energy, comprising compute, switch, and interconnect energy. Given that switch and interconnect components dominate the energy budget, Conv-$\eta$ approaches zero, indicating poor effective utilization. In $\pi^{2}$ architecture, the utilization metric $\pi^{2}$-$\eta$ is defined as the ratio of useful switch and interconnect energy to the total energy consumed by the switch and interconnect. This formulation reflects the integrated nature of computation and communication in $\pi^{2}$ systems, where the switch itself performs both functions efficiently. We distinguish between the portion of switch energy that actively contributes to $\pi^{2}$ operations—denoted $E^{sw}_{comp}$ and comprising delay scheduling, temporal sorting, queue-gating, credit accumulation, and event generation—and the remaining energy that does not participate in computation, including leakage, idle energy, initialization, clock trees, and interface overhead ($E^{sw}_{base}=E_{sw}-E^{sw}_{comp}$) and express $\pi^{2}-\eta$ as $\frac{E_{comp}^{sw}+E_{interconnect}}{E_{sw}+E_{interconnect}}$. The ratio of $E_{comp}^{sw}:E_{sw}$ is projected to be 0.36 in [spower]. The interconnect delays can be characterized and utilized for computation as well. By utilizing these physical effects of interconnects and switch operations for computation rather than treating them as purely communicational overhead, we can push the $\eta$ much closer to 1.
The peak observed in $\pi^{2}-\eta$ in Table 1 arises because, between 2014 and 2018, interconnect energy did not scale at the same rate as switch core and compute energy, causing the interconnect component to dominate the energy budget (45% drop in interconnect energy). By 2022, however, the introduction of co-packaged optics (CPO) significantly reduced interconnect energy (79% drop in interconnect energy), leading to the subsequent drop in $\eta$. It is important to note that the energy-utilization analysis assumes a dedicated-mode deployment, in which the switch fabric is reserved exclusively for $\pi^{2}$ event processing and does not simultaneously forward conventional data traffic.
@@ 4.2 $\pi^{2}$ primitives and standard neural network operations
$\pi^{2}_{K}$ neuron is a variant of the TEMP neuron model: Let $\mathcal{U}=\{t_{1},t_{2},\ldots,t_{d}\}$ denote the set of input event arrival times. For simplicity, we assume that $t_{1}\leq t_{2}\leq\cdots\leq t_{d}$ For a fixed parameter $\gamma\geq 0$, the TEMP neuron generates a spike/event at time T when the following condition is satisfied [temp]:
@@
$\pi^{2}_{K}$ neurons can be interpreted as mimicking inner-product computation on their inputs: At the core of the TEMP neuron model lies margin propagation (MP), a piecewise-linear approximate computing technique introduced in [bias, mulp_mp, temp]. In [mulp_mp], MP was employed to approximate inner products between vectors, and the error characteristics of this approximation were analyzed and quantified. Following a similar reasoning, the TEMP neuron model can also be interpreted as an approximation to the inner product. Let $\mathbf{w},\mathbf{x}\in\mathbb{R}^{D}$ be $D$-dimensional real vectors. For convenience, we define the sets
@@
where $TEMP(\cdot,\gamma)$ denotes the operator defined in Eq. 17. Eq. 20 highlights the differential formulation inherited from MP to approximate the inner product. Since we have established the equivalence between TEMP and the proposed $\pi^{2}_{K}$ neuron, the same relationship can be expressed in terms of $\pi^{2}_{K}$ as
@@ 4.3 Training of $\pi^{2}$ neural networks on GPU
We adopt a knowledge transfer-based training strategy for the $\pi^{2}$-NN, wherein a conventional MAC-based neural network (Teacher network) is first trained to achieve baseline accuracy on a given task. The trained model’s parameters are then used to initialize the $\pi^{2}$-NN, which is subsequently fine-tuned to match the performance of the teacher network. For all software-based simulations, we refer to the conventional MAC-based network as the teacher network and its corresponding $\pi^{2}$-NN implementation as the student network. For training the student network, we employ a distillation loss that combines the conventional cross-entropy loss with a temperature-scaled Kullback-Leibler (KL) divergence between the soft output distributions of the teacher and student networks. At the input layer, to convert the pixel values (X) to differential spike timing (T), as specified in [temp], we use the following strategy.
@@
Training ResNet-9 architecture on the CIFAR-10 dataset: The CIFAR-10 dataset, comprising 50,000 training and 10,000 test images across 10 classes, was preprocessed by applying standard data augmentation techniques commonly used in image classification tasks. Each training image was randomly cropped to $32\times 32$ pixels with a padding of 4 using reflection padding, followed by random horizontal flipping. The pixel values were then normalized using the dataset-specific channel-wise mean and standard deviation values: $(\mu_{R},\mu_{G},\mu_{B})=(0.4914,0.4822,0.4465)$ and $(\sigma_{R},\sigma_{G},\sigma_{B})=(0.2470,0.2435,0.2616)$. The teacher network was trained using the Adam optimizer with a learning rate of $1\times 10^{-2}$, a weight decay of $1\times 10^{-4}$, and a OneCycleLR scheduler to achieve the baseline performance. We initially trained the $\pi^{2}$-NN naively (with the same hyperparameters), from scratch, using backpropagation and achieved a 5% drop from the SOTA accuracy. To improve accuracy, we implemented the KD strategy. To train the $\pi^{2}$-NN student model on CIFAR-10, we adopt a knowledge distillation strategy, initializing the network using a pre-trained MAC-based ResNet9 teacher model. The K values corresponding to the 9 layers of the network were as follows:$K=[16,50,35,35,60,70,80,200,50]$. The $\pi^{2}$ network was trained using the Adam optimizer with a learning rate of $1\times 10^{-2}$, a weight decay of $1\times 10^{-8}$, and a OneCycleLR scheduler for both the experiments. Training was conducted over 60 epochs using a batch size of 16.
Training the ResNet-9 architecture on the CIFAR-100 dataset: We train the teacher ResNet-9 network on the CIFAR-100 dataset, comprising 50,000 training and 10,000 test images across 100 classes. Image preprocessing includes random cropping with 4-pixel padding, horizontal flipping, and normalization using channel-wise mean and standard deviation values of (0.5071, 0.4867, 0.4408) and (0.2675, 0.2565, 0.2761), respectively. The model is trained using the Adam optimizer with a learning rate of $1\times 10^{-2}$ and a weight decay of $1\times 10^{-4}$ for 50 epochs. A OneCycleLR scheduler is employed to modulate the learning rate with a three-phase annealing schedule. The K values used to train the $\pi^{2}$-NN student model on CIFAR-100 were [10, 50, 55, 65, 100, 100, 80, 200, 100]. The network was trained using the Adam optimizer with a learning rate of $1\times 10^{-2}$, a weight decay of $1\times 10^{-8}$, and a OneCycleLR scheduler. Training was conducted over 60 epochs using a batch size of 16. We also trained the $\pi^{2}$-NN naively, from scratch, using backpropagation and achieved a 5% drop from the SOTA accuracy, with the same training setup. To improve accuracy, we implemented the KD strategy.
@@ 4.4 OMNET++ implementation of the MNIST $\pi^{2}$ network
The nodes/neurons are simulated as TSN devices, compatible with the traffic-shaping protocols. The synaptic delays are quantized and encoded within the Priority Code Point (PCP) fields of the incoming Ethernet frames. The OMNeT++ simulator supports a three-bit Priority Code Point (PCP) that is part of the VLAN tag [nesting]. So, the synaptic delays are quantized to 3 bits. The input events are first routed to an Ethernet switch implementing the $\pi^{2}$-ATS protocol. The TET values are computed based on the PCP value of the frame (synaptic delay), and the events are buffered until their TET is reached. The delayed events are then transmitted to another Ethernet switch implementing the $\pi^{2}$-CBS protocol to simulate the $\pi^{2}_{K}$ dynamics. The implementation details are presented in the supplementary section 4.
@@ 4.6 Calibrating for finite-precision, jitter and packet drops
We implement two types of hardware constraints: (i) spiking time jitter and (ii) weight quantization, and (iii) random packet drops. To incorporate weight quantization (Fig. 7A), we employ quantization-aware training (QAT), allowing weights to be represented with reduced bit precision during inference. Specifically, we fine-tune the network for 10 epochs using quantized weights, starting from a pretrained model. To incorporate spiking time jitter (Fig. 7C), a random value of a Gaussian noise of a given standard deviation is added to each spiking time of the TTFS-network inputs and the outputs of each layer, similar to [stanojevic2024, jitter1]. We fine-tune the network for 10 epochs using jitter and quantized weights (8 bits and 4 bits), starting from the pre-trained model. For packet drops (Fig. 7B), we stochastically suppress events within a layer with a preset drop probability and assess the resulting performance; this condition is likewise fine-tuned for 10 epochs from the pretrained initialization. Across all robustness experiments, we use an initial learning rate of 1e-3 and keep all other hyperparameters identical to the original training configuration.
4.7 Training $\pi^{2}_{s}$ networks
We train the $\pi^{2}_{s}$- ResNet-18 network on the CIFAR-10 dataset following a similar strategy of weight porting and knowledge distillation. The model is trained using the Adam optimizer with a learning rate of $1\times 10^{-2}$ and a weight decay of $1\times 10^{-4}$ for 75 epochs. A OneCycleLR scheduler is employed to modulate the learning rate with a three-phase annealing schedule. The sparsity value was set to 1 for the first layer. We experiment with a sparsity value of 1 at the input layer and 0.3 for the remaining layers and report the results. Further details on the training setup and the K values for the 18-layer network can be obtained from our shared code repository.
@@ 4.8 Computational complexity of $\pi^{2}$ networks
where $C_{m}$ and $C_{a}$ represent the computational complexity of performing a multiplication and addition operation. For an n x n $\pi^{2}$ network following the $\pi^{2}_{K}$ neuron paradigm, the number of addition operations performed at every node of the output layer depends on the value of K. The delaying, time sorting, and event dropping operations are inherently implemented using interconnects. The number of operations at the neuron level is
@@
In the hardware implementation of conventional networks, the interconnects do the work of transmitting data to and from memory, external interfaces, communicating modules, and between different submodules within the hardware substrate. The communication complexities and memory accesses are usually ignored while computing the computational complexity in conventional networks [spkdriven]. Accordingly, in both cases, we have only considered the complexity of neuron-level computations. To emphasize the hardware efficiency of this approach, it is noteworthy that a 16-bit full adder typically requires approximately 250–500 transistors [add], while a 16-bit multiplier demands over 4,000 transistors [mul]. Furthermore, the energy cost of a floating-point multiplication can be up to four times that of a floating-point addition [horowitz], underscoring the significant hardware and energy advantages of the $\pi^{2}_{K}$ neuron.
@@ Acknowledgements
The authors would like to thank Haresh Dagale (Principal Research Scientist, Department of Electronic Systems Engineering, IISc) for his insights on communication network protocols. The authors would also like to acknowledge the joint IISc-WashU Memorandum of Understanding, to facilitate the collaboration between the two institutions. M.S.R., C.S.T., and S.C. participated in the Neuromorphic Integrated Circuits workgroup at the Telluride Neuromorphic and Cognition Engineering workshop in 2024, Colorado, USA, which was supported by the National Science Foundation grant ECCS:2332166. M.S.R., C.S.T., and S.C. also acknowledge the 2025 Bangalore Neuromorphic Engineering workshop (BNEW 2025), where some of the $\pi^{2}$-NN discussions occurred. S.C. would like to acknowledge the WashU McDonnell International Scholars Academy for supporting travel to BNEW 2025 at IISc. C.S.T. would like to acknowledge the Pratiksha Trust for funding support.
@@
M.S.R., C.S.T., and S.C. formulated the concept of $\pi^{2}$-NN and the use of interconnect primitives. M.S.R. performed all the simulations and benchmarking. C.B. contributed to improving the training of $\pi^{2}$-NN. All authors/co-authors contributed to the proofreading and writing of the manuscript.
@@
The specific Python codes used for training $\pi^{2}$ networks are available at https://github.com/NeuRonICS-Lab/Processing-in-Interconnect-Codes. Omnet++ codes will be available on request.
@@
Supplementary Information for When Routers, Switches and Interconnects Compute: A Processing-in-Interconnect Paradigm for Scalable Neuromorphic AI
@@ 1 A simple $\pi^{2}$ inspired point-to-point (P2P) processor architecture
Figure S1: A) $\pi^{2}$ inspired P2P architecture. B) Roofline model of the proposed architecture
We propose a point-to-point $\pi^{2}$ network architecture (Fig.S1A) where every pair of neurons is connected through a path with a specific routing delay. The synaptic weights in a conventional neural network correspond to interconnect delays here. The interconnect length—or the specific routing path taken by the data—is determined by the trained synaptic weight magnitudes and is pre-defined. Thus, whenever an input neuron sends out data, it gets delayed automatically before it reaches its destination. Also, the shared interconnect bandwidth to the Rx neuron can only accommodate K events. Since the events arrive in a time-sorted manner, the first K events will inherently be accommodated at the Rx node, and the remaining events will be dropped. This approach effectively integrates processing and memory within the interconnect itself, as it inherently delays, sorts, and drops events, eliminating the need for explicit synaptic weight storage. We assume the neurons possess sufficient local memory and compute to accumulate membrane potential values and generate an event at T (output spike time).
@@
$Ops/s=\frac{Ln^{2}+n+n(Ln-K)}{T}$ (S2)
As can be observed, only increasing the number of interconnects in this architecture, or the bandwidth, can delay, drop, and transmit more events, thus improving the peak throughput of this processor architecture. Additionally, there is no access to any external memory, so the roofline model remains parallel to the X-axis, as shown in Fig. S1B. The latency of computation (T) will likely be further reduced, as the Rx neurons are likely to generate events faster (as they receive more data), which is bound to improve the throughput even further.
@@
If the interconnect between the Tx and Rx nodes has a bandwidth BW of L events/T(s), then the power consumed (W) in T seconds is:
$P_{useful}=\frac{Ln^{2}E_{t}+nE_{c}}{T}$ (S3)
@@
$P_{lost}=\frac{E_{p}n|F-L|_{+}}{T}$ (S4)
where $_{+}=\max(0,.)$ is a ReLU operation. When interconnects contribute to both computation and transmission, the transmission energy becomes an integral part of the system’s operational cost rather than an overhead. The energy/power utilization of the system can be computed to be:
$\eta=1-P_{lost}/(P_{useful}+P_{lost})$ (S5)